Sharon drove 295 miles 5 1/2 hours. she drove at an average rite of 45 mph for the first 3 hours. what was her average rate for the reminder of the trip?
Answer by ankor@dixie-net.com(22740)

Sharon drove 295 miles 5 1/2 hours.
she drove at an average rite of 45 mph for the first 3 hours.
what was her average rate for the reminder of the trip?
:
Let r = the rate for the 2nd part of the trip
5.5 - 3 = 2.5 hrs for the 2nd part
:
Write a distance equation, dist = speed * time
3(45) + 2.5r = 295
135 + 2.5r = 290
2.5r = 295 - 135
2.5r = 160
r = 160/2.5
r = 64 mph on the 2nd part of the trip
:
:
Check the solution by finding the distances
3(45) + 2.5(64) =
135 + 160 = 295